OpenStudy (anonymous):

Calculate module and argument of the following:

OpenStudy (anonymous):

\[e ^{\frac{ 3 }{ 2 }\pi i}\]

OpenStudy (anonymous):

*modulus

OpenStudy (turingtest):

\[\Large e^{i\theta}=\cos\theta+i\sin\theta\]so\[\Large e^{3\pi i\over2}=\cos(\frac{3\pi}2)+i\sin(\frac{3\pi}2)\]what does that simplify to?

OpenStudy (anonymous):

I dont know...